OpenStudy (anonymous):

what is the maximum height to which a motor having a power rating of 20.4 watts can lift a 5.00-kilogram stone vertically in 10.0 seconds?

OpenStudy (kropot72):

Joule = watt-second Energy available from motor = 20.4 * 10 Joules Joule = Newton-meter 20.4 * 10 = 5 * (maximum height) \[maximum\ height=\frac{20.4\times 10}{5}\ meters\]